# Joint Source-Channel Optimization for UAV Video Coding and Transmission

### v1

@@ Abstract
This paper is concerned with unmanned aerial vehicle (UAV) video coding and transmission in scenarios such as emergency rescue and environmental monitoring. Unlike existing methods of modeling video source coding and channel transmission separately, we investigate the joint source-channel optimization issue for video coding and transmission. Particularly, we design eight-dimensional delay-power-rate-distortion models in terms of source coding and channel transmission and characterize the correlation between video coding and transmission, with which a joint source-channel optimization problem is formulated. Its objective is to minimize end-to-end distortion and UAV power consumption by optimizing fine-grained parameters related to UAV video coding and transmission. This problem is confirmed to be a challenging sequential-decision and non-convex optimization problem. We therefore decompose it into a family of repeated optimization problems by Lyapunov optimization and design an approximate convex optimization scheme with provable performance guarantees to tackle these problems. Based on the theoretical transformation, we propose a Lyapunov repeated iteration (LyaRI) algorithm. Extensive experiments are conducted to comprehensively evaluate the performance of LyaRI. Experimental results indicate that compared to its counterparts, LyaRI is robust to initial settings of encoding parameters, and the variance of its achieved encoding bitrate is reduced by 47.74%.
@@ I Introduction
In the application of UAVs, video acquisition and transmission constitute one of their pivotal missions. UAVs are capable of capturing targets through on-board cameras and subsequently transmitting coded and compressed videos via on-board data transmission modules. The prompt acquisition of on-site dynamic information via UAV videos can secure invaluable time for emergency rescue operations. Consequently, the UAV video coding and transmission technology has emerged as a focal point of interest within both academia and industry [4, 5, 6].
@@ I-A State of the Arts
Recent years have witnessed a proliferation of research in the domain of UAV video transmission. The authors in [9] proposed an approach to maximize video utility through the joint optimization of resource allocation and UAV trajectory, thereby reducing operational time of UAV. In [10], the authors introduced a quality-of-experience (QoE)-driven dynamic wireless video broadcasting scheme, aiming to maximize the peak signal-to-noise ratio (PSNR) of reconstructed video by jointly optimizing UAV transmit power and trajectory. The aforementioned studies focused on enhancing throughput of UAV networks and reducing transmission distortion by optimizing UAV trajectories, resource allocation, and transmit power without considering video coding distortion. To tackle this issue, the authors in [11] investigated the optimal deployment of UAVs in UAV-assisted virtual reality (VR) systems with the objective of minimizing video transmission delay while meeting the image resolution requirements. This research focused on resolution, an important indicator of video coding, but did not delve into video coding.
@@
All these studies assumed unlimited encoding time and power consumption. They also focused on the selection of video coding quantization parameters based on the R-D relationship and aimed to achieve the best encoding performance. However, the time and power consumption of video coding have been confirmed to significantly impact the video bitrate and distortion. In [15], the authors explored semi-extreme sparse coding of transform residuals at medium and low bitrates to enhance video compression efficiency. The work in [16] indicated that video bitrate and distortion were primarily determined by the distribution characteristics of the transform residuals. Transform residuals were mainly influenced by motion estimation (ME) and quantization parameters. The accuracy of ME was closely related to the sum of absolute differences (SAD) of macroblocks. The number of calculations of macroblock SAD was proportional to the encoding complexity. Considering that the time and power consumption of video coding were increasing functions of encoding complexity, the authors in [16] deduced a quantitative relationship between video coding delay, power, rate, and distortion and extended the traditional R-D model to the d-P-R-D model. Based on this model, a new rate control (RC) method was proposed, aiming to minimize encoding distortion under constraints of rate, delay, and power. However, there are two aspects for improvement in this research. Firstly, the work is conducted for the earlier video coding standard advanced video coding (AVC). High efficiency video coding (HEVC) and versatile video coding (VVC), as the subsequent video coding standards[17], offer lower coding bitrates under the same video quality conditions. Therefore, it is necessary to further investigate the d-P-R-D model for the HEVC standard. Secondly, although the end-to-end delay composition of video transmission systems was analyzed in [16], the proposed d-P-R-D model was only for video source coding. To achieve better end-to-end video transmission QoE, it is essential to design a joint source-channel video coding and transmission scheme.
@@
From the perspective of joint source-channel design, efficient video source coding can reduce the video bitrate while maintaining the same video quality. The reduction in video bitrate leads to decreasing transmit power consumption and transmission delay. However, efficient video source coding implies higher computational complexity, which result in increased power consumption and delay in video coding. Given the constraints of end-to-end latency, increasing encoding latency to achieve better compression performance may reduce the transmission latency budget, subsequently increasing the risk of distortion during transmission. Similarly, changes in encoding power consumption can affect the power consumption of video transmission under a given power constraint of video transmission system. Researchers have conducted extensive study on joint source-channel optimization [18, 19, 20, 21, 22, 23, 24, 25]. For instance, the work in [18] proposed a new scheme for dynamically adjusting the bitrate based on channel signal-to-noise ratio (SNR) and image content. In [19], the authors presented a joint source-channel adaptive RC scheme for wireless image transmission based on image feature mapping, entropy awareness, and channel conditions. The authors in [21] introduced a novel video codec structure based on compressed sensing. They also investigated the minimization of system power consumption by jointly optimizing video coding bitrate and multi-path transmission bitrate under distortion and energy consumption constraints. The work in [23] investigated the trade-off between latency, power, and distortion and achieved deep cross-layer optimization of video coding, queuing control, and wireless transmission through a joint lossy compression and power allocation scheme. However, the aforementioned research focused on the field of wireless video transmission and did not specifically target UAV video transmission. To address this, the authors in [8] conducted cross-layer design research on UAV-based video streaming transmission. They proposed a quality of service (QoS) strategy that dynamically adjusted the transmission mode according to network load, latency, and packet loss rate. Nevertheless, they focused on the overall system design of multi-link parallel transmission, without delving into the issue of UAV video coding. Considering the dynamic nature and power constraints of UAVs, it is necessary to conduct joint resource allocation and control for UAV video coding and transmission under the constraints of end-to-end latency and power consumption. In this way, the QoE of video transmission at the receiving end can be improved.
@@ I-B Motivation and Contributions
Overall, the optimization of UAV location/trajectory and network resource allocation to enhance the performance of UAV video transmission has emerged as a hot research topic. However, the issue of joint source-channel coding control and resource optimization allocation for UAV video coding transmission has not been adequately explored. This paper proposes a cross-layer d-P-R-D joint design scheme based on HEVC coding. The objective is to minimize end-to-end distortion and power consumption by optimizing video coding parameters and the allocation of UAV transmit power under the constraints of end-to-end latency, power budget, and source-channel rate matching. The main research content and contributions of this paper are summarized as follows:
@@
4) Finally, we design extensive experiments to quantitatively analyze the stability of the proposed algorithm and the effectiveness of the joint source-channel optimization mechanism. We thoroughly discuss the impact of various parameters on the performance of the algorithm. Experimental results demonstrate that the proposed algorithm achieves mean-rate stability. Compared to benchmark algorithms, the proposed algorithm is less affected by initial settings of encoding parameters, and the variance of its achieved encoding bitrate is reduced by 47.74%.
@@
Figure 1: A UAV video coding and transmission scene.
@@ II-A AtG Channel Model
This paper primarily considers the application scenarios where UAVs are utilized for emergency search and rescue, geological disaster monitoring, forest fire detection, and other missions that involve capturing and transmitting videos in a timely manner, as illustrated in Fig. 1. The scenario addressed in this paper encompasses an emergency communication vehicle (ECV) and a UAV. Given the potential for road traffic disruptions in the areas under surveillance, such as forests or post-disaster zones, the ECV is deployed in proximity to the affected region to launch the UAV. The UAV is tasked with conducting surveillance and reconnaissance missions, capturing video footage for transmission. The captured video is encoded and compressed by the UAV and then promptly delivered to the ECV. To facilitate the construction of a mathematical model for the video transmission process, the temporal domain is discretized into a sequence of time slots, denoted by $t=\{1,2,\ldots\}$. Due to the limited coverage range of the UAV, without loss of generality, this paper assumes that the UAV flies along a circular trajectory to extend the coverage range. The two-dimensional (2-D) horizontal coordinates of the UAV at time slot $t$ can be represented as $\bm{q_{uav}}\left(t\right)={[{x_{uav}}\left(t\right),{y_{uav}}\left(t\right)]^{\rm{T}}}$, and the position of the ECV is denoted as $\bm{q_{ecv}}={[{x_{ecv}},{y_{ecv}}]^{\rm{T}}}$. Consider that the UAV flies at a constant altitude $H$, and the altitude of the ECV is negligible compared to $H$. The horizontal projection distance between the UAV and the ECV at time slot can be expressed as $dis(t)=\sqrt{{{({x_{uav}}\left(t\right)-{x_{user}})}^{2}}+{{({y_{ecv}}\left(t\right)-{y_{ecv}})}^{2}}}$.
The UAV and the ECV communicate through an AtG link. In the AtG link, the UAV and the ECV typically establish line of sight (LoS) communication with a certain probability. This probability is contingent upon various factors including the flying altitude of the UAV, the horizontal projection distance between the UAV and the ECV, and the flying environment. The commonly used expression for the probability of LoS communication is[26].
@@
where $a$ and $b$ are constant coefficients related to the flying environment of the UAV, and $\theta(t)=\frac{{180}}{\pi}\arctan(\frac{H}{{dis(t)}})$ represents the elevation angle between the ECV and the UAV. Disregarding the altitude of the ECV and the height of the UAV’s antenna, the transmission loss of the AtG link can be expressed as follows[26]
@@ II-B1 Video Bitrate and Distortion Models
In the IPPPP coding mode, both encoding bitrate and distortion of P-frames in block-based hybrid video coding can be approximately represented as functions of the standard deviation of the transform residual and the quantization step $Q$ [16]. Based on the research findings in literature[27], the functional expression $\sigma$ can be obtained through nonlinear regression analysis. This paper selects standard video test sequences CITY (CIF, 352*288) and Coastguard (CIF, 352*288) and employs the HM-16.20 encoder to obtain encoding data samples. In the low-latency configuration file encoder_lowdelay_P_main.cfg of HM, since quantization parameter $QP$, search range $\lambda$, and reference frames $\chi$ are independent parameters, their individual impacts on $\sigma$ can be assessed separately. It should be noted that there is a one-to-one mapping relationship between the quantization parameter $QP$ mentioned in HM-16.20 and the quantization step $Q$ involved in the model. This paper first fixes $QP$ and $\chi$, and changes the value of $\lambda$ to obtain transform residual data and calculate $\sigma$. Through nonlinear regression analysis, it has been found that $\sigma$ is approximately exponentially related to $\lambda$. Secondly, by fixing $\lambda$ and $\chi$, and changing $QP$, it can be concluded that $\sigma$ is approximately linearly related to $QP$. Additionally, it is observed that the rate of change of $\sigma$ versus $\chi$ is much smaller than that of $\sigma$ versus $QP$, as well as $\sigma$ versus $\lambda$. It indicates that $\chi$ has a relatively smaller impact on $\sigma$ compared to $QP$ and $\lambda$. To reduce the computational complexity, this paper assigns a given value ${\chi_{0}}=1$ to $\chi$[16]. Therefore, this paper adopts the following expression to calculate $\sigma$
@@
When auxiliary information such as macroblock types is ignored, the bitrate of video coding can be approximated as the entropy of the transform-quantized coefficients[28]. There are various assumptions for the distribution of transform residuals. For instance, the research in [29] assumed that transform residuals followed a zero-mean generalized Gaussian distribution (GGD). However, the GGD has some limitations in practicality due to its multiple control parameters. In AVC, the Cauchy distribution has also been utilized, but the convergence issue of its mean and variance must be considered when employing the Cauchy distribution. The Laplace distribution achieves a good balance between computational complexity and model accuracy and has been widely applied [28, 30]. This paper therefore chooses the Laplacian distribution as the statistical model for transform residuals.
@@
For source coding, where the transform residuals follow a zero-mean Laplacian independent and identically distributed (i.i.d.) model, the relationship between the video coding bitrate ${R_{e}}(\lambda,Q)$ and $Q$ as well as $\lambda$ can be mathematically represented as
@@
It should be noted that, to model the relationship between video coding bitrate and time slot, the symbol ${R_{e}}(\lambda,Q;t)$ is also utilized in the subsequent discussions to denote the video coding bitrate.
Compared to AVC, HEVC employs different RC strategies[31]. In the research of HEVC, a hyperbolic model is commonly utilized to describe the relationship between video coding bitrate $R$ and distortion $D$[32], i.e. $D(R)=C\cdot{R^{-K}}$, where, $C$ and $K$ are parameters related to video content. After obtaining the video coding bitrate through source entropy, this paper models the relationship between video coding distortion and bitrate as follows

### v2

@@ Abstract
This paper is concerned with unmanned aerial vehicle (UAV) video coding and transmission in scenarios such as emergency rescue and environmental monitoring. Unlike existing methods of modeling UAV video source coding and channel transmission separately, we investigate the joint source-channel optimization issue for video coding and transmission. Particularly, we design eight-dimensional delay-power-rate-distortion models in terms of source coding and channel transmission and characterize the correlation between video coding and transmission, with which a joint source-channel optimization problem is formulated. Its objective is to minimize end-to-end distortion and UAV power consumption by optimizing fine-grained parameters related to UAV video coding and transmission. This problem is confirmed to be a challenging sequential-decision and non-convex optimization problem. We therefore decompose it into a family of repeated optimization problems by Lyapunov optimization and design an approximate convex optimization scheme with provable performance guarantees to tackle these problems. Based on the theoretical transformation, we propose a Lyapunov repeated iteration (LyaRI) algorithm. Both objective and subjective experiments are conducted to comprehensively evaluate the performance of LyaRI. The results indicate that, compared to its counterparts, LyaRI achieves better video quality and stability performance, with a 47.74% reduction in the variance of the obtained encoding bitrate.
@@ I Introduction
In the application of UAVs, video acquisition and transmission constitute one of their pivotal missions. UAVs are capable of capturing targets through on-board cameras and subsequently transmitting coded and compressed videos via on-board data transmission modules. The prompt acquisition of on-site dynamic information via UAV videos can secure invaluable time for emergency rescue operations. As a result, the UAV video coding and transmission technology has emerged as a focal point of interest within both academia and industry [4, 5, 6].
@@ I-A State of the Arts
Recent years have witnessed a proliferation of research in the domain of UAV video transmission. The authors in [9] proposed an approach to maximize video utility through the joint optimization of resource allocation and UAV trajectory, thereby reducing operational time of the UAV. In [10], the authors introduced a quality of experience (QoE) driven dynamic wireless video broadcasting scheme, aiming to maximize the peak signal-to-noise ratio (PSNR) of reconstructed video by jointly optimizing UAV transmit power and trajectory. The aforementioned studies focused on enhancing throughput of UAV networks and reducing transmission distortion by optimizing UAV trajectories, resource allocation, and transmit power without considering video coding distortion. To tackle this issue, the authors in [11] investigated the optimal deployment of UAVs in UAV-assisted virtual reality (VR) systems with the objective of minimizing video transmission delay while meeting the image resolution requirements. This research focused on resolution, an important indicator of video coding, but did not delve into video coding.
@@
All these studies assumed unlimited encoding time and power consumption. They also focused on the selection of video coding quantization parameters based on the R-D relationship and aimed to achieve the best encoding performance. However, the time and power consumption of video coding have been confirmed to significantly impact the video bitrate and distortion. In [15], the authors explored semi-extreme sparse coding of transformed residuals at medium and low bitrates to enhance video compression efficiency. The work in [16] indicated that video bitrate and distortion were primarily determined by the distribution characteristics of the transformed residuals. Transformed residuals were mainly influenced by motion estimation (ME) and quantization parameters. The accuracy of ME was closely related to the sum of absolute differences (SAD) of macroblocks. The number of calculations of macroblock SAD was proportional to the encoding complexity. Considering that the time and power consumption of video coding were increasing functions of encoding complexity, the authors in [16] deduced a quantitative relationship between video coding delay, power, rate, and distortion, and extended the traditional R-D model to a delay-power-rate-distortion (d-P-R-D) model. Based on this model, a new rate control (RC) method was proposed, aiming to minimize encoding distortion under constraints of rate, delay, and power. However, there are two aspects for improvement in this research. Firstly, the work is conducted for the earlier video coding standard advanced video coding (AVC). High efficiency video coding (HEVC) and versatile video coding (VVC), as the subsequent video coding standards[17], offer lower coding bitrates under the same video quality conditions. Therefore, it is necessary to further investigate the d-P-R-D model for the HEVC standard. Secondly, although the end-to-end delay composition of video transmission systems was analyzed in [16], the proposed d-P-R-D model was only for video source coding. To achieve better end-to-end video transmission QoE, it is essential to design a joint source-channel video coding and transmission scheme.
@@
From the perspective of joint source-channel design, efficient video source coding can reduce the video bitrate while maintaining the same video quality. The reduction in video bitrate leads to decreasing transmit power consumption and transmission delay. However, efficient video source coding implies higher computational complexity, which result in increased power consumption and delay in video coding. Given the constraints of end-to-end latency, increasing encoding latency to achieve better compression performance may reduce the transmission latency budget, subsequently increasing the risk of distortion during transmission. Similarly, changes in encoding power consumption can affect the power consumption of video transmission under a given power constraint of video transmission system. Researchers have extensively studied joint source-channel optimization [18, 19, 20, 21, 22, 23, 24, 25]. For instance, the work in [18] proposed a new scheme for dynamically adjusting the bitrate based on channel signal-to-noise ratio (SNR) and image content. In [19], the authors presented a joint source-channel adaptive RC scheme for wireless image transmission based on image feature mapping, entropy awareness, and channel conditions. The authors in [21] introduced a novel video codec structure based on compressed sensing. They also investigated the minimization of system power consumption by jointly optimizing video coding bitrate and multi-path transmission bitrate under distortion and energy consumption constraints. The work in [23] investigated the trade-off between latency, power, and distortion and achieved deep cross-layer optimization of video coding, queuing control, and wireless transmission through a joint lossy compression and power allocation scheme. However, the aforementioned research focused on the field of wireless video transmission and did not specifically target UAV video transmission. To address this, the authors in [8] conducted cross-layer design research on UAV-based video streaming transmission. They proposed a quality of service (QoS) strategy that dynamically adjusted the transmission mode according to network load, latency, and packet loss rate. Nevertheless, they focused on the overall system design of multi-link parallel transmission, without delving into the issue of UAV video coding. Considering the dynamic nature and power constraints of UAVs, it is necessary to conduct joint resource allocation and control for UAV video coding and transmission under the constraints of end-to-end latency and power consumption. In this way, the QoE of video transmission at the receiving end can be improved.
@@ I-B Motivation and Contributions
Overall, the optimization of UAV location/trajectory and network resource allocation to enhance the performance of UAV video transmission has emerged as a hot research topic. However, the issue of joint source-channel coding control and resource optimization allocation for UAV video coding and transmission has not been adequately investigated. This paper proposes a joint source-channel d-P-R-D design scheme based on HEVC coding. The objective is to minimize end-to-end distortion and power consumption by optimizing video coding parameters and the allocation of UAV transmit power under the constraints of end-to-end latency, power budget, and source-channel rate matching. The main research content and contributions of this paper are summarized as follows:
@@
4) Finally, we design objective and subjective experiments to quantitatively analyze the stability of the proposed algorithm and the effectiveness of the joint source-channel optimization mechanism. We thoroughly discuss the impact of various parameters on the performance of the algorithm. Experimental results demonstrate that the proposed algorithm achieves mean-rate stability. Compared to benchmark algorithms, the proposed algorithm achieves better video quality and stability performance, with a 47.74% reduction in the variance of the obtained encoding bitrate.
@@
Figure 1: A UAV video coding and transmission scenario.
@@ II-A AtG Channel Model
This paper primarily considers the application scenarios where UAVs are deployed for emergency search and rescue, geological disaster monitoring, forest fire detection, and other missions that involve encoding and transmitting videos in a timely manner. Particularly, the scenario investigated in this paper encompasses a UAV and an emergency communication vehicle (ECV). Given the potential for road traffic disruptions in areas such as mountainous and post-disaster zones, the ECV is deployed in proximity to the affected region to launch the UAV. The UAV is tasked with conducting surveillance and reconnaissance missions, capturing video footage for transmission, as illustrated in Fig. 1. The captured video is encoded and compressed by the UAV and then promptly delivered to the ECV. To facilitate the construction of a mathematical model for the video transmission process, the temporal domain is discretized into a sequence of time slots, denoted by $t=\{1,2,\ldots\}$. Due to its limited surveillance coverage range, the UAV is deployed to fly along a circular trajectory to extend the coverage range. The two-dimensional (2-D) horizontal coordinate of the UAV at time slot $t$ is represented as $\bm{q_{uav}}\left(t\right)={[{x_{uav}}\left(t\right),{y_{uav}}\left(t\right)]^{\rm{T}}}$, and the position of the ECV is denoted as $\bm{q_{ecv}}={[{x_{ecv}},{y_{ecv}}]^{\rm{T}}}$. Consider that the UAV flies at a constant altitude $H$, and the altitude of the ECV is negligible compared to $H$. The horizontal projection distance between the UAV and the ECV at time slot $t$ can be expressed as $dis(t)=\sqrt{{{({x_{uav}}\left(t\right)-{x_{user}})}^{2}}+{{({y_{ecv}}\left(t\right)-{y_{ecv}})}^{2}}}$.
The UAV and the ECV communicate through an AtG link. In the AtG link, the UAV and the ECV typically establish line of sight (LoS) communication with a certain probability. This probability is contingent upon various factors including the flying altitude of the UAV, the horizontal projection distance between the UAV and the ECV, and the flying environment[26]. The commonly used expression for the probability of LoS communication is as follows
@@
where $a$ and $b$ are constant coefficients related to the flying environment of the UAV, and $\theta(t)=\frac{{180}}{\pi}\arctan(\frac{H}{{dis(t)}})$ represents the elevation angle between the ECV and the UAV.
Disregarding the altitude of the ECV and the height of the UAV’s antenna, the transmission loss of the AtG link [26] can be expressed as follows
@@ II-B1 Video Bitrate and Distortion Models
In the IPPPP coding mode, both encoding bitrate and distortion of P-frames in block-based hybrid video coding can be approximately represented as functions of the standard deviation of transformed residuals and the quantization step $Q$ [16]. Based on the research findings in literature[27], the functional expression $\sigma$ can be obtained through nonlinear regression analysis. This paper selects standard video test sequences CITY (CIF, 352*288) and Coastguard (CIF, 352*288) and employs the HM-16.20 encoder to obtain encoding data samples. In the low-latency configuration file encoder_lowdelay_P_main.cfg of HM, since quantization parameter $QP$, search range $\lambda$, and reference frames $\chi$ are independent parameters, their individual impacts on $\sigma$ can be assessed separately. It should be noted that there is a one-to-one mapping relationship between the quantization parameter $QP$ mentioned in HM-16.20 and the quantization step $Q$ involved in the model. This paper first fixes $QP$ and $\chi$, and changes the value of $\lambda$ to obtain transformed residual data and calculate $\sigma$. Through nonlinear regression analysis, it has been found that $\sigma$ is approximately exponentially related to $\lambda$. Secondly, by fixing $\lambda$ and $\chi$, and changing $QP$, it can be concluded that $\sigma$ is approximately linearly related to $QP$. Additionally, it is observed that the rate of change of $\sigma$ versus $\chi$ is much smaller than that of $\sigma$ versus $QP$, as well as $\sigma$ versus $\lambda$. It indicates that $\chi$ has a relatively smaller impact on $\sigma$ compared to $QP$ and $\lambda$. To reduce the computational complexity, this paper assigns a given value ${\chi_{0}}=1$ to $\chi$[16]. Therefore, this paper adopts the following expression to calculate $\sigma$
@@
When auxiliary information such as macroblock types is ignored, the bitrate of video coding can be approximated as the entropy of the transform-quantized coefficients[28]. There are various assumptions for the distribution of transformed residuals. For instance, the research in [29] assumed that transformed residuals followed a zero-mean generalized Gaussian distribution (GGD). However, the GGD has some limitations in practicality due to its multiple control parameters. In AVC, the Cauchy distribution has also been utilized, but the convergence issue of its mean and variance must be considered when employing the Cauchy distribution. The Laplace distribution achieves a good balance between computational complexity and model accuracy and has been widely applied [28, 30]. This paper therefore chooses the Laplacian distribution as the statistical model for transformed residuals.
@@
For video coding, where the transformed residuals follow a zero-mean Laplacian independent and identically distributed (i.i.d.) model, the relationship between the video coding bitrate ${R_{e}}(\lambda,Q)$ and $Q$ as well as $\lambda$ can be mathematically represented as
@@
It should be noted that, to model the relationship between video coding bitrate and time slot, the symbol ${R_{e}}(\lambda,Q;t)$ is also utilized in the subsequent discussions to denote the video coding bitrate. Compared to AVC, HEVC employs different RC strategies[31]. In the research of HEVC, a hyperbolic model is commonly utilized to describe the relationship between video coding bitrate $R$ and distortion $D$[32], i.e. $D(R)=C\cdot{R^{-K}}$, where, $C$ and $K$ are parameters related to video content. After obtaining the video coding bitrate through source entropy, this paper models the relationship between video coding distortion and bitrate as follows